# Semiclassical measure of the propagation between two topological insulators

### v2

@@ 1.3 Main result
then, there exist a vanishing sequence ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ of positive numbers and a family of measurable maps ${t\mapsto\left(\mathrm{d}\gamma^{t}_{n}\right)_{n\in\mathbb{Z}}}$ valued in the set of scalar non negative Radon measures on $\mathbb{R}^{2}$, such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\int_{\mathbb{R}}\Xi(t)\left\langle a\psi^{\varepsilon_{k}}_{t},\psi^{\varepsilon_{k}}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\mathrm{d}t\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}}\int_{\mathbb{R}}\Xi(t)\mathrm{tr}\Big{(}a\big{(}\gamma(s)\big{)}\mathrm{d}\rho_{t}(s)\Big{)}\mathrm{d}t$
with for all $s\in\mathbb{R}$,
$\mathrm{d}\rho_{t}(s)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}}\begin{pmatrix}1&\frac{\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\mathrm{i}\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\gamma^{t}_{n}(s,\mathrm{d}\sigma),$
where for all $n\in\mathbb{Z}$, $\mathrm{d}\gamma_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\mathrm{d}\gamma_{n}^{t}&=&\partial_{\sigma}\mathrm{d}\gamma_{n}^{t}\ \partial_{s}\lambda_{n}-\partial_{s}\mathrm{d}\gamma_{n}^{t}\ \partial_{\sigma}\lambda_{n}\\ \mathrm{d}\gamma_{n}^{t}|_{t=0}&=&\mathrm{d}\gamma_{n,0}\end{array}\right.,$ (6)
and $\mathrm{d}\gamma_{n,0}$ is determined by the initial condition.
We will see in Section 3.3 a more general result (Theorem 3.10), with initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$.
Remark (Link between Theorem 1.1 and Theorem 1.4).
With the notation of Theorem 1.4, the measure $\mathrm{d}\gamma_{0}^{t}$ is the measure associated to the part $\mathrm{WP}^{\varepsilon}_{x_{t},0}$ of the solution of Theorem 1.1. Likewise, the measures $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}_{\neq 0}}$ are associated to the $\mathrm{L}^{\infty}$-remainder part of the solution of Theorem 1.1. Notably, if for all $n\in\mathbb{Z}_{\neq 0},\mathrm{d}\gamma_{n,0}=0$ and $\mathrm{d}\gamma_{0,0}=1$, then for all $(t,s,\sigma)\in\mathbb{R}^{3}$ and for all $n\in\mathbb{Z}_{\neq 0}$,
$\mathrm{d}\gamma_{0}^{t}(s,\sigma)=\delta(s-t)\delta(\sigma)\quad,\quad\mathrm{d}\gamma_{n}^{t}(s,\sigma)=0,$
where $\delta$ is the Dirac mass over 00.
Theorem 1.4 gives a description of the evolution of the position density when initial data are wave packets. But Theorem 3.10 allow us to consider any initial data in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ concentrating near the curve $\mathcal{C}$ to describe their asymptotic position density behavior. A consequence of this theorem is the following one.
Theorem 1.5.
Assume Assumption 1.2 and 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ a vanishing sequence such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
@@
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition, then,
@@
$\mathrm{d}\nu\mathds{1}_{\mathrm{E}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{d}\gamma^{t}_{n}\right)\mathds{1}_{\mathrm{E}}\mathrm{d}t$
@@
where $\left(\mathrm{d}\gamma_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (6).
This theorem gives a lower bound estimation of the position density at time $t$ of the evolution of any initial data concentrating over the curve $\mathrm{E}$.
1.4 Main ideas of the proofs and organisation of the article
There are two main ideas in the proof of Theorem 3.10.
The first step is to transform the operator $\mathrm{H}_{\varepsilon}$ into a model operator as if the curve $\mathrm{E}$ were exactly a line. Let us discuss the case of the line.
If the interface is a line defined by a mass $\mathrm{m}$ such that for all $x\coloneqq(x_{1},x_{2})\in\mathbb{R}^{2},\mathrm{m}(x)=\widetilde{m}(x_{1})x_{2}$, then the operator $\mathrm{H}_{\varepsilon}$ is unitary equivalent to the following one,
This operator is $\sqrt{\varepsilon}$-semiclassical with respect to $x_{1}$ only. Thus, it will be considered as the quantization of an operator-valued symbol. It leads us to the second idea of the proof.
The second step is to build a two-scale quantization that fits the scale of our new operator. In fact, our analysis relies on a microlocal approach, replacing densities $\left(\left|\psi_{t}^{\varepsilon}(x)\right|^{2}\mathrm{d}x\right)_{\varepsilon>0}$ by the Wigner transform [26, 15]. The Wigner transform provides us with a description of the dynamics for a very general class of equations and initial data, allowing us to deal with quantum observables. Therefore, we will study the quadratic quantity (3), with an observable $a$ depending of position and momentum, i.e.
$a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi},\mathbb{C}^{2,2}\right)\longmapsto\left\langle a\left(\cdot{},\varepsilon\mathrm{D}_{x}\right)\psi^{\varepsilon}_{t},\psi^{\varepsilon}_{t}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
We recall basic facts about these notions in Section 2. However, to precisely analyze what happens near the curve $\mathcal{C}$, we have to take into account how $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ concentrates on the curve $\mathcal{C}$ with respect to the scale $\sqrt{\varepsilon}$. We introduce in Section 3 the two-scale Wigner measure that be of crucial use for proving our main results Theorems 3.8 and 3.10. Finally, we prove Theorems 1.41.53.8 and 3.10 in Section 4.
1.5 Notation
We denote by $\mathbb{S}^{2}$ the unit sphere of $\mathbb{R}^{3}$.
We denote by $v\otimes w$ the $2$ by $2$ matrix $v\,^{t}w$ for $(v,w)\in\left(\mathbb{C}^{2}\right)^{2}$.
We denote by ${\rm tr}$ the trace in $\mathcal{L}\left(\mathbb{C}^{2}\right)$ and by $\mathrm{Tr}$ the trace in $\mathcal{L}\left(\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)\right)$.
Let $\mathcal{H}$ a separable Hilbert space and $n$ an integer, for $f,g\in\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote the usual inner product by
$\left\langle f,g\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)}\coloneqq\int_{\mathbb{R}^{n}}\left\langle f(x),g(x)\right\rangle_{\mathcal{H}}\,\mathrm{d}x.$
For $A$ an operator acting on $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$, we denote by $A^{\dagger}$ the formal adjoint operator of $A$ associated to the usual scalar product of $\mathrm{L}^{2}\left(\mathbb{R}^{n},\mathcal{H}\right)$.
We define the semiclassical Fourier transform for all $\varepsilon>0$ and for $f\in\mathrm{L}^{1}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$ by
and we denote by $\widehat{f}$ the Fourier transform of $f$ defined by $\widehat{f}\coloneqq\mathcal{F}_{1}f$.
2 Semiclassical analysis
@@
Our problem requires an analysis in the phase space. We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1) and its weak limits in the space of distributions that are called Wigner measures. We recall the definition and the main properties of the Wigner transform in Section 2.1 below. Then, we apply these results to the solution of our problem in Section 2.2.
2.1 Wigner transform and Wigner measure
Let ${(f^{\varepsilon})_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. The Wigner transform of ${(f^{\varepsilon})_{\varepsilon>0}}$ is defined as follow, for all $(x,\xi)\in\mathbb{R}^{2}\times\mathbb{R}^{2}$,
$\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\coloneqq\int_{\mathbb{R}^{2}}e^{i\xi\cdot y}f^{\varepsilon}\left(x+\frac{\varepsilon y}{2}\right)\otimes\overline{f^{\varepsilon}\left(x-\frac{\varepsilon y}{2}\right)}\frac{\mathrm{d}y}{(2\pi)^{2}}.$
It was introduced by Wigner in [25] and in [26]. One can check that $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is valued in the set of Hermitian matrices and $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]\in\mathrm{L}^{2}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\cap\mathscr{C}^{0}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$. Moreover, one recovers formally the density in position through the relation
$\int_{\mathbb{R}^{2}}{\rm tr}\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\mathrm{d}\xi=\left|f^{\varepsilon}(x)\right|_{\mathbb{C}^{2}}^{2}.$
The function $\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right]$ is usually considered as a distribution on the phase space $\mathbb{R}^{4}=\mathbb{R}^{2}_{x}\times\mathbb{R}^{2}_{\xi}$. Moreover, we have for all $a\in\mathscr{C}^{\infty}_{c}(\mathbb{R}^{4},\mathbb{C}^{2,2})$,
$\mathrm{I}^{\varepsilon}_{f}(a)\coloneqq\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{W}^{\varepsilon}\left[f^{\varepsilon}\right](x,\xi)\Big{)}\mathrm{d}x\mathrm{d}\xi=\left\langle\mathrm{Op}_{\varepsilon}(a)f^{\varepsilon},f^{\varepsilon}\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)},$
where $\mathrm{Op}_{\varepsilon}(a)$ is the semiclassical pseudo-differential operator of symbol $a$ obtained by the Weyl quantization
$\mathrm{Op}_{\varepsilon}(a)f(x)\coloneqq\int_{\mathbb{R}^{4}}a\left(\frac{x+y}{2},\xi\right)f(y)e^{\mathrm{i}\frac{\xi\cdot(x-y)}{\varepsilon}}\frac{\mathrm{d}y\mathrm{d}\xi}{(2\pi\varepsilon)^{2}},\quad f\in\mathscr{S}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),\ x\in\mathbb{R}^{2}.$
The Calderón-Vaillancourt Theorem [4] asserts that the family of operators $\left(\mathrm{Op}_{\varepsilon}(a)\right)_{\varepsilon>0}$ is uniformly bounded with respect to $\varepsilon$ in $\mathcal{L}\left(L^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)\right)$: there exists $(C,N)\in\mathbb{R}_{>0}\times\mathbb{N}_{\geqslant 0}$ such that for all $a=(a_{i,j})_{1\leqslant i,j\leqslant 2}\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right),$
$\left\lVert\mathrm{Op}_{\varepsilon}(a)f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\leqslant C\,\max_{1\leqslant i,j\leqslant 2}\,\max_{\left|\alpha\right|+\left|\beta\right|\leqslant N}\left\lVert\partial^{\alpha}_{x}\partial^{\beta}_{\xi}a_{i,j}\right\rVert_{\mathrm{L}^{\infty}\left(\mathbb{R}^{4}\right)}\left\lVert f\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$ (7)
The Wigner transform is not a positive distribution. However, its weak limits are positive measures. This is due to the weak Gårding inequality. Adapting [11] Proposition 2.13 to the matrix case, for all $a\in\mathscr{C}^{\infty}_{c}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$ non negative, for all $\delta>0$, there exists $C_{\delta}>0$, for all $f\in\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$,
$\left\langle\mathrm{Op}_{\varepsilon}(a)f,f\right\rangle_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}\geqslant-\left(\delta+C_{\delta}\varepsilon\right)\left\lVert f\right\rVert^{2}_{\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)}.$
There exists a subsequence $\left(f_{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$ and a positive matrix of Radon measures $\mathrm{d}\mu=(\mathrm{d}\mu_{i,j})_{1\leqslant i,j\leqslant 2}$ such that for all $a\in\mathscr{S}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)$,
$\mathrm{I}^{\varepsilon_{n}}_{f}(a)\underset{n\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{4}}{\rm tr}\Big{(}a(x,\xi)\mathrm{d}\mu(x,\xi)\Big{)}.$ (8)
The measure $\mathrm{d}\mu$ is called the Wigner measure or the semiclassical measure of $\left(f^{\varepsilon_{n}}\right)_{n\in\mathbb{N}}$. A proof of existence is recalled in Section B.1 for families valued in a separable Hilbert space.
Example 2.1.
@@ 2.1 Wigner transform and Wigner measure
Now, we apply these first properties on solutions to equation (1).
2.2 Application to the Dirac equation
In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1) with the initial normalized condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. We revisit in this paragraph the notions introduced in the preceding one for this special family.
2.2.1 $\varepsilon$-oscillation of solution to Dirac equation
Lemma 2.2 (Conservation of $\varepsilon$-oscillating).

### v3

@@ 3.3 Proof of applications
then, applying Corollary 1.7 and Corollary 2.7, there exist a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ and a two-scale semiclassical measure at finite distance ${M^{t}\mathrm{d}\nu^{t}}$, such that for all observables $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{4},\mathbb{C}^{2,2}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
$\mathcal{I}^{\varepsilon_{k}}_{\psi}(\Xi,a)\underset{k\to+\infty}{\longrightarrow}\frac{1}{2}\sum_{n\in\mathbb{Z}}\int_{\mathbb{R}^{3}}\Xi(t)\mathrm{tr}\left(a\big(\mathbf{t}(s),0\big)\begin{pmatrix}1&\frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{-\mathrm{i}\theta(s)}\\ \frac{-\sigma}{\lambda_{n}(s,\sigma)}e^{\mathrm{i}\theta(s)}&1\end{pmatrix}\mathrm{d}\nu^{t}_{n}(s,\sigma)\right)\mathrm{d}t,$
with, for all $n\in\mathbb{Z}$, $\nu_{n}^{t}$ satisfies the following equation
$\left\{\begin{array}[]{rll}\partial_{t}\ \nu_{n}^{t}&=&\left\{\nu_{n}^{t},\lambda_{n}\right\}\\ \nu_{n}^{t}|_{t=0}&=&\nu_{n,0}\end{array}\right$.,
with $\nu_{n,0}=\mathrm{tr}\Big(\Pi_{n}M^{0}\Big)\nu^{0}$ and $M^{0}\mathrm{d}\nu^{0}$ is a two-scale semiclassical measure associated with the initial data.
We conclude thanks to the conservation of $\varepsilon$-oscillation (Lemma A.2). ∎
Proof of Corollary 1.9.
Assume Assumption 1.1 and Assumption 1.3. Let ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ a uniformly bounded family of $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$, $\nu$ a non-negative scalar Radon measure on $\mathbb{R}_{t}\times\mathbb{R}_{x}^{2}$ and a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k\in\mathbb{N}}}$ such that for all $(a,\Xi)\in\mathscr{C}_{c}^{\infty}\left(\Omega,\mathbb{C}\right)\times\mathscr{C}_{c}^{\infty}(\mathbb{R},\mathbb{C})$,
@@
where ${\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}}$ solves (1.1) with ${\left(\psi_{0}^{\varepsilon}\right)_{\varepsilon>0}}$ as initial condition. A consequence of Theorem 1.5 and Remark 2.6 is
@@
$\mathrm{d}\nu\mathds{1}_{\mathcal{C}}\geqslant\sum_{n\in\mathbb{Z}}\Phi^{*}\left(\mathrm{Tr}\Big(\Pi_{n}(s,\sigma)\Big)\mathrm{d}\nu^{t}_{n}\right)\mathds{1}_{\mathcal{C}}\mathrm{d}t$
@@
where $\left(\nu_{n}^{t}\right)_{n\in\mathbb{Z}}$ is defined by (1.20). We conclude with $\mathrm{Tr}\left(\Pi_{n}(s,\sigma)\right)=1$. ∎
Appendix A Semiclassical analysis
@@
Our problem requires an analysis in phase space. We consider the Wigner transform of the solution $\left(\psi_{t}^{\varepsilon}\right)_{\varepsilon>0}$ of equation (1.1) and its weak limits in the space of distributions that are called Wigner measures. We apply the main properties of the Wigner transform (introduced in Section 1.2) to the solution of our problem in Section A.1.
First, let us briefly introduce an important example and some definitions.
Example A.1.
@@
Now, we apply these first properties on solutions to equation (1.1).
A.1 Application to Dirac equation
In this section, we consider $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ the solution to equation (1.1) with the normalized initial condition $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$. We revisit in this paragraph the notions introduced in the preceding one for this special family.
Lemma A.2 (Conservation of $\varepsilon$-oscillating).